OpenStudy (anonymous):

A student's body temperature is 98.6 F.What is this temperature in Celsius scale?

OpenStudy (anonymous):

Are you good with formulas?

OpenStudy (anonymous):

If so, then I will give you the formula to solve that.

OpenStudy (anonymous):

36.8 degrees celcius

OpenStudy (anonymous):

Here is the formula: F=( C+40 ) (9/5)-40= 37 degrees celsius. Here is how you do that: 98.6 + 40 = 138.6 138.6 divided by 9 = 15.4 15.4 times 5 = 77 77-40 = 37 There you go! :)

OpenStudy (anonymous):

Thank you